OpenStudy (anonymous):

MEDAL AND FOLLOW 1 QUESTION

OpenStudy (anonymous):

Mario Agnello is opening a new pizza shop in town. He plans to offer a twelve-inch diameter pizza and an eighteen-inch diameter pizza. He has set the price of a twelve-inch cheese pizza at eight dollars, based on the amount of ingredients needed to cover that size crust. If Mario would like to keep the pricing proportional, what should he charge for an eighteen-inch cheese pizza? Show and explain your solution.

OpenStudy (anonymous):

do you know how to find the area of the pizzas

OpenStudy (anonymous):

nooo

OpenStudy (anonymous):

well the equation is pi(r^2) and d=2r

OpenStudy (anonymous):

so for the 12-inch pizza the area is 36pi and for the 18 inch pizza the area would be 81pi do you understand?

OpenStudy (anonymous):

81s the answer?

OpenStudy (anonymous):

no you need to know the areas so you can make a proportion

OpenStudy (anonymous):

soo to make the proportion you would have 8/36pi=x/81pi because 8 is the price of the 12inch pizza and 36pi is the area of the twelve inch pizza and it is x over 81pi because you are trying to find the price of the 18inch pizza. do you understand how to solve this equation?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok so tell me the answer when you figure it out

OpenStudy (anonymous):

its confusing

OpenStudy (anonymous):

8/36pi=x/81pi so you cross multiply to get 36pix=648pi then you would divide each side by 36pi x=18 so the price would be $18

OpenStudy (anonymous):

you understand?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

any other questions

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

The area of triangle A B C is twelvesquare centimeters. Each of the dimensions of triangle D E F is 3 times the dimension of triangle A B C. What is the area of triangle D E F? Triangle A B C and triangle D E F are placed side-by-side. They are both similar triangles. Triangle A B C is smaller than triangle D E F. 4 square centimeters thirty-six square centimeters seventy-two square centimeters one hundred eight square centimeters

OpenStudy (anonymous):



OpenStudy (anonymous):

sorry I cant help you here im more of and algebra person then geometry. Ill have to think back two years ago. I might be able to find it if you give me some time. Sorry again

OpenStudy (anonymous):

I think the answer might be the second on but im not sure

OpenStudy (anonymous):

the answer is the first one because the ratio of the triangle is 3:1 and instead of them trying to get you to square every dimension which would make triangle DEF 9 cm bigger they are just want you to double the base and the height which would be 2 dimensions squared which would be the same as saying2^2:1 as the ratio or 4cm bigger then triangle ABC.

OpenStudy (anonymous):

its askin what the area of def if @taylorjohnson287